OpenStudy (anonymous):

solve 0.6x+3+3≤1.4x-2

jimthompson5910 (jim_thompson5910):

0.6x+3+3<=1.4x-2 10(0.6x+3+3)<=10(1.4x-2) 6x+30+30<=14x-20 6x+60<=14x-20 6x<=14x-20-60 6x-14x<=-20-60 -8x<=-20-60 -8x<=-80 x>=(-80)/(-8) x>=10